# Comparison results for Markov tree distributions

### v1

@@ Appendix B Proofs of Section 1
$\begin{split}E[\phi(Z^{(k-1)})]&=E[\varphi^{(0)}(Z^{(k-1)}_{i},Z^{(k-1)}_{d_{i}})]\\ &=E[\varphi^{(1)}(Z^{(k-1)}_{i},Z^{(k-1)}_{c_{i}},Z^{(k-1)}_{d_{j_{2}}},\ldots,Z^{(k-1)}_{d_{j_{m}}})]\\ &=\cdots\\ &=E[\varphi^{(m-1)}(Z^{(k-1)}_{i},Z^{(k-1)}_{c_{i}},Z^{(k-1)}_{d_{j_{m}}})]\\ &=E[\varphi^{(m)}(Z^{(k-1)}_{i},Z^{(k-1)}_{c_{i}})],\end{split}$ (68)
which shows (64). The proof for Equation (65) follows similarly replacing $k-1$ by $k$ in (68). Hence, we obtain from (64) and (65) that
$E[\phi(Z^{(k-1)})]$ $=E[\varphi^{(m)}(Z^{(k-1)}_{i},Z^{(k-1)}_{c_{i}})]$
$=E[\varphi^{(m)}(X_{i},X_{c_{i}})]$
$\leq E[\varphi^{(m)}(Y_{i},Y_{c_{i}})]$
$=E[\varphi^{(m)}(Z^{(k)}_{i},Z^{(k)}_{c_{i}})]$
$=E[\phi(Z^{(k)})],$
For the second and third equality, we use that $(Z^{(k-1)}_{i},Z^{(k-1)}_{c_{i}})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(X_{i},X_{c_{i}})$ and $(Z^{(k)}_{i},Z^{(k)}_{c_{i}})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(Y_{i},Y_{c_{i}})\,,$ which follows from the definition of $Z^{(k-1)}$ and $Z^{(k)}$, respectively. For the inequality, we apply Lemma 1.2 using on the one hand that, by Markov tree dependence, the random variables $X_{c_{i}}=(X_{j_{1}},\ldots,X_{j_{m}})$ are conditionally independent given $X_{i}$ and the random variables $Y_{c_{i}}=(Y_{j_{1}},\ldots,Y_{j_{m}})$ are conditionally independent given $Y_{i}$. On the other hand, $X_{j_{\ell}}\uparrow_{st}X_{i}$ by assumption (i), $Y_{j_{\ell}}\uparrow_{st}Y_{i}$ by assumption (ii), and $(X_{i},X_{j_{\ell}})\leq_{sm}(Y_{i},Y_{j_{\ell}})$ for $\ell\in\{1,\ldots,m\}\,.$
@@
The comparison result formulated in Proposition B.5 is an extension of Lemma 1.2 to tree structures. However, it does not generalize Proposition 1.1 since the SI assumptions in Proposition B.5 are more stringent when applied to chains. It turns out that these SI assumptions on $Y$ can be omitted in one direction along a specified path $P$ such that an generalization of Proposition 1.1 for trees can be formulated. This result is presented in Theorem 1.3 which can be proven by employing Proposition B.5 and an alternative adjustment of the dependencies, as described for the proof of Proposition B.5. As we show in the following proof of Theorem 1.3, we can omit the SI assumption on $X$ for one of the edges connected with the root and we can omit the SI assumptions on $Y$ along a path $P$ through the tree starting with the root node and terminating at a leaf node whenever it has finite length. The proof of Theorem 1.3 is presented in the following.
  • Proof of Theorem 1.3
    Similar to the proof of Proposition B.5, we can assume that $d+1\coloneqq|N|<\infty$. We show that $X\leq_{sm}Y$. The proof for $X\geq_{sm}Y$ follows analogously. Denote by $F=(F_{n})_{n\in N}$ for $F_{n}=F_{X_{n}}=F_{Y_{n}}$ the marginal distribution functions of $X$ and $Y$. Moreover, for $e=(i,j)\in E$, denote by $B_{e}$ and $C_{e}$ any copula for $(X_{i},X_{j})$ and $(Y_{i},Y_{j})$, respectively, and define $B=(B_{e})_{e\in E}$ and $C=(C_{e})_{e\in E}$. Then $X\sim\mathcal{M}(F,B,T)$ and $Y\sim\mathcal{M}(F,C,T)$, see Proposition 2.5. We enumerate the path $P=p(0,\ell]=:\{\ell_{1},\dots,\ell_{m}\}$ such that $(\ell_{j},\ell_{j+1})\in E$ for all $j=0,\dots,m-1$, where $\ell_{0}:=0$ and $\ell_{m}=\ell\,.$
@@
    Recall that $k^{*}$ is due to the assumptions of Theorem 1.3 a child node of the root 00 and lies in $P$ if and only if $\deg(0)=1\,.$ Note also that, for a node $i\,,$ the set $d_{i}$ is defined as the set of descendants of $i\,.$ Each set of nodes $N_{k}$ forms a tree $T_{k}=(N_{k},E_{k})$ along with its corresponding set of edges $E_{k}$ defined by
@@
    i.e., $\{N_{k}\}_{k=0,\ldots,m}$ is a partition of $N$ and $\{E_{k}\}_{k=0,\ldots,m}$ is a partition of the set of edges $E$ without the edges that are contained in the specified path $P$ and without the edge between the root node and $k^{*}\,.$ In the upcoming steps, we will sucessively replace the bivariate copulas of $X$ with those of $Y$. For $e\in E$, consider the sequences $A^{1}=(A^{1}_{e})_{e\in E}$ and $A^{2}=(A^{2}_{e})_{e\in E}$ of bivariate copulas given by
    $A_{e}^{1}$$\coloneqq\begin{cases}C_{e}&\text{if }e=(0,k^{*}),\\ B_{e}&\text{else, }\end{cases}$
    $A_{e}^{2}$$\coloneqq\begin{cases}C_{e}&\text{if }e\in\{(0,k^{*})\}\cup E_{0}\\ B_{e}&\text{else. }\end{cases}$
@@
    $D^{0}_{e}$$\coloneqq\begin{cases}A_{e}^{2}&\text{if }\operatorname{deg}(0)>1,\\ B_{e}&\text{if }\operatorname{deg}(0)=1,\\ \end{cases}$
    and for $k=1,\dots,m$ the sequences $D^{2k-1}=(D^{2k-1}_{e})_{e\in E}$ and $D^{k}=(D^{k}_{e})_{e\in E}$ of bivariate copulas by
@@
    Note that by definition $D^{2m}=C$. Keep in mind that our aim is to change all dependencies described by $(B_{e})_{e\in E}$ with the dependencies described by $(C_{e})_{e\in E}$. The difference between the copula sets $D^{2k-1}$ and $D^{2k}$ lies in the replacement of the copula $B_{(\ell_{k-1},\ell_{k})}$ by $C_{(\ell_{k-1},\ell_{k})}$. On the other hand, the difference between $D^{2k}$ and $D^{2k+1}$ is given by the replacement of the set of copulas $(B_{e})_{e\in E_{k+1}}$ by $(C_{e})_{e\in E_{k+1}}$. Let
@@
    be random vectors on $\mathbb{R}^{d+1}$. The following part of the proof is divided into two steps. In the first step we show that $X\leq_{sm}Z^{(0)}$ and in the second step we show that $Z^{(0)}\leq_{sm}Y$.
@@
    is supermodular. Note that we do not need any SI assumption on $(X_{0},X_{k^{*}})$. By the the disintegration theorem and the Markov property, we have that
@@
    where the second equality in the first and second line follow from $W^{(1)}_{N_{0}}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{N_{0}}$ and $W^{(2)}_{N_{0}}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{N_{0}}\,$, respectively. Assumptions (i)- (iii), imply that the subvectors $(X_{n})_{n\in N_{0}}$ and $(Y_{n})_{n\in N_{0}}$ fulfill the assumptions of Proposition B.5. Thus, (75) implies $W^{(1)}\leq_{sm}W^{(2)}$, which proves (71) using $W^{(2)}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Z^{(0)}$ by the definition of $W^{(2)}$ and $Z^{(0)}\,.$
@@
    The proof of (76) is devided into two parts. First, we show that
@@
    To prove (77) let $\phi\colon\mathbb{R}^{d+1}\to\mathbb{R}$ be a bounded and supermodular function and $k\in\{1,\dots,m\}$. If $|N_{k}|\leq 1$ we have that $Z^{(2k-2)}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Z^{(2k-1)}$ and thus $Z^{(2k-2)}\leq_{sm}Z^{(2k-1)}$. So assume that $|N_{k}|>1\,.$ Since, by assumption (ii), $Y_{j}\uparrow_{st}Y_{i}$ for all $(i,j)\in\bigcup_{u=0}^{k-1}E_{u}$ and $Y_{k^{*}}\uparrow_{st}Y_{0}$ if $\deg(0)>1$ and $Y_{i}\uparrow_{st}Y_{j}$ for all $(i,j)\in\bigcup_{u=1}^{k-1}\{(\ell_{u-1},\ell_{u})\}$ we have that $Z^{(2k-2)}_{j}\uparrow_{st}Z^{(2k-2)}_{i}$ for all $(i,j)\in\bigcup_{u=0}^{k-1}E_{u}$ and $Z^{(2k-2)}_{k^{*}}\uparrow_{st}Z^{(2k-2)}_{0}$ if $\deg(0)>1$ and $Z^{(2k-2)}_{i}\uparrow_{st}Z^{(2k-2)}_{j}$ for all $(i,j)\in\bigcup_{u=1}^{k-1}\{(\ell_{u-1},\ell_{u})\}$ Further, by assumption (i), $X_{j}\uparrow_{st}X_{i}$ for all $(i,j)\in\bigcup_{u=k+1}^{m}E_{u}\cup\bigcup_{u=k}^{m}\{(\ell_{u-1},\ell_{u})\}$ and consequently $Z^{(2k-2)}_{j}\uparrow_{st}Z^{(2k-2)}_{i}$ for all $(i,j)\in\bigcup_{u=k+1}^{m}E_{u}\cup\bigcup_{u=k}^{m}\{(\ell_{u-1},\ell_{u})\}$. Note that if $k=1$ and $\operatorname{deg}(0)=1$, it holds that $|N_{1}|=\{0\}$, which belongs to the previous case. Now, we obtain from Lemma B.4 (ii), applied on the subtree with nodes $(N\backslash N_{k})\cup\{\ell_{k-1}\}$ and root $\ell_{k-1}$ by setting $K=N_{k}\backslash\{\ell_{k-1}\}$ and $J=N\backslash N_{k}$, that the function $\varphi^{(2k-1)}$ given by
    $\varphi^{(2k-1)}(x_{N_{k}})\coloneqq\int\phi(x)P^{Z^{(2k-2)}_{N\backslash N_{k}}|Z^{(2k-2)}_{\ell_{k-1}}=x_{\ell_{k-1}}}(dx_{N\backslash N_{k}})$ (79)
@@
    where the first equality in each line follows by the disintegration theorem and the Markov property and the second ones since $X_{N_{k}}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Z_{N_{k}}^{(2k-2)}$ and $Y_{N_{k}}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Z_{N_{k}}^{(2k-1)}$ by definition of $Z^{(2k-2)}$ and $Z^{(2k-1)}$ in (70). Assumptions (i)- (iii) of Theorem 1.3, imply that the subvectors $(X_{n})_{n\in N_{k}}$ and $(Y_{n})_{n\in N_{k}}$ fulfill the assumptions of Proposition B.5 which implies $E[\varphi^{(2k-1)}(X_{N_{k}})]\leq E[\varphi^{(2k-1)}(Y_{N_{k}})]$. Together with (80), we obtain (77).
    To prove (78), let again $\phi\colon\mathbb{R}^{d+1}\to\mathbb{R}$ be a bounded and supermodular function and $k\in\{1,\dots,m\}$. Since, by assumption (ii), $Y_{j}\uparrow_{st}Y_{i}$ for all $(i,j)\in\bigcup_{u=0}^{k}E_{u}$ and $Y_{k^{*}}\uparrow_{st}Y_{0}$ if $\deg(0)>1$ and $Y_{i}\uparrow_{st}Y_{j}$ for all $(i,j)\in\bigcup_{u=1}^{k-1}\{(\ell_{u-1},\ell_{u})\}$, we have that $Z^{(2k-1)}_{j}\uparrow_{st}Z^{(2k-1)}_{i}$ for all $(i,j)\in\bigcup_{u=0}^{k}E_{u}$ and $Z^{(2k-1)}_{k^{*}}\uparrow_{st}Z^{(2k-1)}_{0}$ if $\deg(0)>1$ and $Z^{(2k-1)}_{i}\uparrow_{st}Z^{(2k-1)}_{j}$ for all $(i,j)\in\bigcup_{u=1}^{k-1}\{(\ell_{u-1},\ell_{u})\}$. Further, by assumption (i), $X_{j}\uparrow_{st}X_{i}$ for all $(i,j)\in\bigcup_{u=k+1}^{m}E_{u}\cup\bigcup_{u=k+1}^{m}\{(\ell_{u-1},\ell_{u})\}$ and consequently $Z^{(2k-1)}_{j}\uparrow_{st}Z^{(2k-1)}_{i}$ for all $(i,j)\in\bigcup_{u=k+1}^{m}E_{u}\cup\bigcup_{u=k+1}^{m}\{(\ell_{u-1},\ell_{u})\}$. We apply Lemma B.4 (ii) twice and find that the function $\varphi^{(2k)}$ given by
    $\begin{split}&\varphi^{(2k)}(x_{\ell_{k-1}},x_{\ell_{k}})\\ &\coloneqq\int\int\phi(x)P^{Z^{(2k-1)}_{N\backslash(\{\ell_{k-1},\ell_{k}\}\cup d_{\ell_{k}})}|Z^{(2k-1)}_{\ell_{k-1}}=x_{\ell_{k-1}}}(dx_{N\backslash(\{\ell_{k-1},\ell_{k}\}\cup d_{\ell_{k}})})P^{Z^{(2k-1)}_{d_{\ell_{k}}}|Z^{(2k-1)}_{\ell_{k}}=x_{\ell_{k}}}(dx_{d_{\ell_{k}}})\end{split}$ (81)
@@
    To show positive supermodular dependence of $X$, we apply the first statement on $X$ and $Y:=X^{\perp}$. Note that the random vector $X^{\perp}$ fulfills assumption (ii) trivially. Since $X$ fulfills assumption (i), we have for $(i,j)\in E\backslash\{(0,k^{*})\}$ that $X_{j}\uparrow_{st}X_{i}$ which implies $(X_{i},X_{j})\geq_{sm}(X^{\perp}_{i},X^{\perp}_{j})=(Y_{i},Y_{j})$, see (14). Hence, with the additional assumption of positive supermodular dependence of $(X_{0},X_{k^{*}})$, we obtain $(X_{i},X_{j})\geq_{sm}(Y_{i},Y_{j})$ for all $(i,j)\in E\,,$ i.e., assumption (iii) holds for all $(i,j)\in E$. From the first part, we now conclude that $X\geq_{sm}Y=X^{\perp}$.
    A similar argument yields positive supermodular dependence of $Y$ under assumption (ii) and positive supermodular dependence of $(Y_{i},Y_{j})$ for $j\in P\cap L\,.$blacksquare
Appendix C Proofs of Section 2
  • Proof of Proposition 2.5
    For $|N|<\infty$, we prove the statement by an induction over the number of nodes $N=\{0,\dots,d\}$. Starting with $|N|=2$, the statement is trivial. For the induction step, assume that the statement is true for each tree $T$ with nodes $N=\{0,\dots,n\}$ for a fixed $n\in\{1,\ldots,d-1\}\,.$ We need to show the statement for $N=\{0,\dots,n+1\}$. By re-enumeration we may assume that the node $n+1$ is a leaf in $N$ with parent node $n$, i.e., $c_{n+1}=\emptyset$ and $p_{n+1}=n$, see Definition B.1. We define the subtree $T^{\prime}=(N^{\prime},E^{\prime})$ with nodes $N^{\prime}=\{0,\dots,n\}$ and edges $E^{\prime}=\{(j,k)\in E\mid j,k\neq n+1\}$, which is a tree with $|N^{\prime}|-1=n$ nodes. By the induction hypothesis, there is a Markov tree distribution $\mu_{\mathcal{T}^{\prime}}$ specified through $((F_{j})_{j\in N^{\prime}},T^{\prime},(B_{e})_{e\in E^{\prime}})$. Let $(X_{n},X_{n+1})$ be a bivariate random vector having distribution function $F_{X_{n},X_{n+1}}:=B_{n,n+1}(F_{n},F_{n+1})$ defined through Sklar’s theorem. For Borel sets $A\in\mathscr{B}(\mathbb{R}^{n+1})$, $B\in\mathscr{B}(\mathbb{R})$ we define the distribution $\mu_{\mathcal{T}}$ on $\mathcal{B}(\mathbb{R}^{n+2})$ by

### v3

@@ B.1 Proofs of Section 1
[…]
which shows (61). The proof for Equation (62) follows similarly replacing k1k-1 by $k$ in (65). Hence, we obtain from (61) and (62) that
$E[\phi(Z^{(k-1)})]$ $=E[\varphi^{(m)}(Z^{(k-1)}_{i},Z^{(k-1)}_{c_{i}})]=E[\varphi^{(m)}(X_{i},X_{c_{i}})]$
$\leq E[\varphi^{(m)}(Y_{i},Y_{c_{i}})]=E[\varphi^{(m)}(Z^{(k)}_{i},Z^{(k)}_{c_{i}})]=E[\phi(Z^{(k)})],$
For the second and third equality, we use that $(Z^{(k-1)}_{i},Z^{(k-1)}_{c_{i}})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(X_{i},X_{c_{i}})$ and $(Z^{(k)}_{i},Z^{(k)}_{c_{i}})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(Y_{i},Y_{c_{i}}),$ which follows from the definition of $Z^{(k-1)}$ and $Z^{(k)}$. For the inequality, we apply Lemma 1.2 using on the one hand that, by Markov tree dependence, the random variables $X_{c_{i}}=(X_{j_{1}},\ldots,X_{j_{m}})$ are conditionally independent given $X_{i}$, and the random variables $Y_{c_{i}}=(Y_{j_{1}},\ldots,Y_{j_{m}})$ are conditionally independent given $Y_{i}$. On the other hand, $X_{j_{\ell}}\uparrow_{st}X_{i}$ by assumption (i), $Y_{j_{\ell}}\uparrow_{st}Y_{i}$ by assumption (ii), and $(X_{i},X_{j_{\ell}})\leq_{sm}(Y_{i},Y_{j_{\ell}})$ for $\ell\in\{1,\ldots,m\}.$
@@
The comparison result formulated in Proposition B.5 extends Lemma 1.2 to tree structures. However, it does not generalize Proposition 1.1 as the SI assumptions in Proposition B.5 are more stringent when applied to chains. In the following proof of Theorem 1.3, it turns out that these SI assumptions on $Y$ can be omitted in one direction along a specified path $P$ starting at the root node and—provided the path has finite length—terminating at a leaf node. Consequently, Theorem 1.3 also extends Proposition 1.1 to tree structures. Additionally, we can omit the SI assumption on $X$ for one of the edges connected with the root.
Similar to the proof of Proposition B.5, we can assume that $d+1\coloneqq|N|<\infty$. We show that $X\leq_{sm}Y$. The proof for $X\geq_{sm}Y$ follows analogously. Denote by $F=(F_{n})_{n\in N}$ for $F_{n}=F_{X_{n}}=F_{Y_{n}}$ the marginal distribution functions of $X$ and $Y$. Moreover, for $e=(i,j)\in E$, denote by $B_{e}$ and $C_{e}$ any copula for $(X_{i},X_{j})$ and $(Y_{i},Y_{j})$, respectively, and define $B=(B_{e})_{e\in E}$ and $C=(C_{e})_{e\in E}$. Then $X\sim\mathcal{M}(F,B,T)$ and $Y\sim\mathcal{M}(F,C,T)$; see Proposition 2.4. We enumerate the path $P=p(0,\ell]=:\{\ell_{1},\dots,\ell_{m}\}$ such that $(\ell_{j},\ell_{j+1})\in E$ for all $j=0,\dots,m-1$, where $\ell_{0}:=0$ and $\ell_{m}=\ell.$
@@
Recall that $k^{*}$ is due to the assumptions of Theorem 1.3 a child node of the root 0 and lies in $P$ if and only if $\deg(0)=1.$ Note also that, for a node $i,$ the set $d_{i}$ is defined as the set of descendants of $i.$ Each set of nodes $N_{k}$ forms a tree $T_{k}=(N_{k},E_{k})$ along with its corresponding set of edges $E_{k}$ defined by
@@
i.e., $\{N_{k}\}_{k=0,\ldots,m}$ is a partition of $N$ and $\{E_{k}\}_{k=0,\ldots,m}$ is a partition of the set of edges $E$ without the edges that are contained in the specified path $P$ and without the edge between the root node and $k^{*}.$ In the upcoming steps, we will successively replace the bivariate copulas of $X$ with those of $Y$. For $e\in E$, consider the sequences $A^{1}=(A^{1}_{e})_{e\in E}$ and $A^{2}=(A^{2}_{e})_{e\in E}$ of bivariate copulas given by
$A_{e}^{1}$ $\coloneqq\begin{cases}C_{e}&\text{if }e=(0,k^{*}),\\ B_{e}&\text{else, }\end{cases}\quad\text{and}\quad A_{e}^{2}\coloneqq\begin{cases}C_{e}&\text{if }e\in\{(0,k^{*})\}\cup E_{0}\\ B_{e}&\text{else. }\end{cases}$
@@
[…]
and for $k=1,\dots,m$ the sequences $D^{2k-1}=(D^{2k-1}_{e})_{e\in E}$ and $D^{2k}=(D^{2k}_{e})_{e\in E}$ of bivariate copulas by
@@
Note that, by definition, $D^{2m}=C$. Keep in mind that our aim is to replace all dependencies described by $(B_{e})_{e\in E}$ by the dependencies described by $(C_{e})_{e\in E}$. The difference between the copula sets $D^{2k-1}$ and $D^{2k}$ lies in the replacement of the copula $B_{(\ell_{k-1},\ell_{k})}$ by $C_{(\ell_{k-1},\ell_{k})}$. On the other hand, the difference between $D^{2k}$ and $D^{2k+1}$ is given by the replacement of the set of copulas $(B_{e})_{e\in E_{k+1}}$ by $(C_{e})_{e\in E_{k+1}}$. Let
@@
be random vectors on $\mathbb{R}^{d+1}$. The following part of the proof is divided into two steps. In the first step we show that $X\leq_{sm}Z^{(0)}$, and in the second step we show that $Z^{(0)}\leq_{sm}Y$.
@@
is supermodular. Note that we do not need any SI assumption on $(X_{0},X_{k^{*}})$. By the disintegration theorem and the Markov property, we have that
@@
where the second equality in the first and second line follow from $W^{(1)}_{N_{0}}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{N_{0}}$ and $W^{(2)}_{N_{0}}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{N_{0}}\,$, respectively. Assumptions (i) – (iii), imply that the subvectors $(X_{n})_{n\in N_{0}}$ and $(Y_{n})_{n\in N_{0}}$ fulfill the assumptions of Proposition B.5. Thus, (72) implies $W^{(1)}\leq_{sm}W^{(2)}$, which proves (68) using $W^{(2)}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Z^{(0)}$ by the definition of $W^{(2)}$ and $Z^{(0)}.$
@@
The proof of (73) is divided into two parts. First, we show that
@@
[…]
@@
where the first equality in each line follows by the disintegration theorem and the Markov property and the second ones since $X_{N_{k}}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Z_{N_{k}}^{(2k-2)}$ and $Y_{N_{k}}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Z_{N_{k}}^{(2k-1)}$ by definition of $Z^{(2k-2)}$ and $Z^{(2k-1)}$ in (67). Assumptions (i) – (iii) of Theorem 1.3, imply that the subvectors $(X_{n})_{n\in N_{k}}$ and $(Y_{n})_{n\in N_{k}}$ fulfill the assumptions of Proposition B.5 which implies $E[\varphi^{(2k-1)}(X_{N_{k}})]\leq E[\varphi^{(2k-1)}(Y_{N_{k}})]$. Together with (77), we obtain (74).
To prove (75), let again $\phi\colon\mathbb{R}^{d+1}\to\mathbb{R}$ be a bounded and supermodular function and $k\in\{1,\dots,m\}$. Since, by assumption (ii), $Y_{j}\uparrow_{st}Y_{i}$ for all $(i,j)\in\bigcup_{u=0}^{k}E_{u}$, and $Y_{k^{*}}\uparrow_{st}Y_{0}$ if $\deg(0)>1$, and $Y_{i}\uparrow_{st}Y_{j}$ f.a. $(i,j)\in\bigcup_{u=1}^{k-1}\{(\ell_{u-1},\ell_{u})\}$, we have that $Z^{(2k-1)}_{j}\uparrow_{st}Z^{(2k-1)}_{i}$ for all $(i,j)\in\bigcup_{u=0}^{k}E_{u}$ and $Z^{(2k-1)}_{k^{*}}\uparrow_{st}Z^{(2k-1)}_{0}$ if $\deg(0)>1$ and $Z^{(2k-1)}_{i}\uparrow_{st}Z^{(2k-1)}_{j}$ for all $(i,j)\in\bigcup_{u=1}^{k-1}\{(\ell_{u-1},\ell_{u})\}$. Further, by assumption (i), $X_{j}\uparrow_{st}X_{i}$ for all $(i,j)\in\bigcup_{u=k+1}^{m}E_{u}\cup\bigcup_{u=k+1}^{m}\{(\ell_{u-1},\ell_{u})\}$ and, thus, $Z^{(2k-1)}_{j}\uparrow_{st}Z^{(2k-1)}_{i}$ for all $(i,j)\in\bigcup_{u=k+1}^{m}E_{u}\cup\bigcup_{u=k+1}^{m}\{(\ell_{u-1},\ell_{u})\}$. We apply Lemma B.4 (ii) twice and find that the function $\varphi^{(2k)}$ given by
$\begin{split}&\varphi^{(2k)}(x_{\ell_{k-1}},x_{\ell_{k}})\\ &\coloneqq\int\int\phi(x)P^{Z^{(2k-1)}_{N\backslash(\{\ell_{k-1},\ell_{k}\}\cup d_{\ell_{k}})}|Z^{(2k-1)}_{\ell_{k-1}}=x_{\ell_{k-1}}}(dx_{N\backslash(\{\ell_{k-1},\ell_{k}\}\cup d_{\ell_{k}})})\\ &\hskip 170.71652ptP^{Z^{(2k-1)}_{d_{\ell_{k}}}|Z^{(2k-1)}_{\ell_{k}}=x_{\ell_{k}}}(dx_{d_{\ell_{k}}})\end{split}$ (78)
@@
To show positive supermodular dependence of $X$, we apply the first statement on $X$ and $Y:=X^{\perp}$. Note that the random vector $X^{\perp}$ fulfills assumption (ii) trivially. Since $X$ fulfills assumption (i), we have for $(i,j)\in E\backslash\{(0,k^{*})\}$ that $X_{j}\uparrow_{st}X_{i}$, which implies $(X_{i},X_{j})\geq_{sm}(X^{\perp}_{i},X^{\perp}_{j})=(Y_{i},Y_{j})$; see (8). Hence, with the additional assumption of positive supermodular dependence of $(X_{0},X_{k^{*}})$, we obtain $(X_{i},X_{j})\geq_{sm}(Y_{i},Y_{j})$ for all $(i,j)\in E,$ i.e., assumption (iii) holds for all $(i,j)\in E$. From the first part, we now conclude that $X\geq_{sm}Y=X^{\perp}$.
A similar argument yields positive supermodular dependence of $Y$ under assumption (ii) and positive supermodular dependence of $(Y_{i},Y_{j})$ for $j\in P\cap L.$ $\blacksquare$
Since the supermodular order is a pure dependence order, it requires the marginal distributions to coincide. Hence, from condition (iii) of Theorem 1.3 (i.e., $(X_{i},X_{j})\leq_{sm}(Y_{i},Y_{j})$ for all $(i,j)\in E$) it follows that
$F_{X_{n}}=F_{Y_{n}}=:F_{n}\quad\text{for all }n\in N.$
Further, as a consequence of (6), there exist bivariate copulas for $(X_{i},X_{j})$ and $(Y_{i},Y_{j})$ that are pointwise ordered, i.e.,
$C_{X_{i},X_{j}}\leq_{lo}C_{Y_{i},Y_{j}}\quad\text{for all }e=(i,j)\in E.$
Condition (i) of Theorem 1.3 (i.e., $X_{j}\uparrow_{st}X_{i}$ for $(i,j)\in E$, $j\neq k^{*}$), means that the bivariate random vector $(X_{i},X_{j})$ is CIS. Since, for bivariate random vectors, the SI property is invariant under increasing transformations, it follows that there exists a bivariate copula $B_{e}$ such that
$B_{e}\text{ is SI}\quad\text{and}\quad B_{e}=C_{X_{i},X_{j}}\text{ on }\mathsf{Ran}(F_{i})\times\mathsf{Ran}(F_{j})\quad\text{for }e=(i,j)\in E,\,j\neq k^{*}.$
Let $D^{\top}$ denote the transpose of a bivariate copula $D,$ defined as $D^{\top}(u,v):=D(v,u)$ for all $(u,v)\in[0,1]^{2}.$ Then condition (ii) of Theorem 1.3 (i.e., $Y_{i}\uparrow_{st}Y_{j}$ if $j\notin L$, and $Y_{j}\uparrow_{st}Y_{i}$ if $j\notin P$, $e=(i,j)\in E$) means that there exists a bivariate copula $C_{e}$ such that
[…]
Finally, we have decomposed the distributions of $X$ and $Y$ into bivariate tree specifications with identical marginal distribution functions $F=(F_{n})_{n\in N}$, and with pointwise ordered SI and CI copulas $B=(B_{e})_{e\in E}$ and $C=(C_{e})_{e\in E}$. Due to Proposition 2.4, every bivariate tree specification yields a uniquely determined Markov tree realization, which proves the results incorporating some additional SI conditions. $\blacksquare$
B.2 Proofs of Section 2
For $|N|<\infty$, we prove the statement by an induction over the number of nodes $N=\{0,\dots,d\}$. Starting with $|N|=2$, the statement is trivial. For the induction step, assume that the statement is true for each tree $T$ with nodes $N=\{0,\dots,n\}$ for a fixed $n\in\{1,\ldots,d-1\}.$ We need to show the statement for $N=\{0,\dots,n+1\}$. By re-enumeration we may assume that the node $n+1$ is a leaf in $N$ with parent node $n$, i.e., $c_{n+1}=\emptyset$ and $p_{n+1}=n$; see Definition B.1. We define the subtree $T^{\prime}=(N^{\prime},E^{\prime})$ with nodes $N^{\prime}=\{0,\dots,n\}$ and edges $E^{\prime}=\{(j,k)\in E\mid j,k\neq n+1\}$, which is a tree with $|N^{\prime}|-1=n$ nodes. By the induction hypothesis, there is a Markov tree distribution $\mu_{\mathcal{T}^{\prime}}$ specified through $((F_{j})_{j\in N^{\prime}},T^{\prime},(B_{e})_{e\in E^{\prime}})$. Let $(X_{n},X_{n+1})$ be a bivariate random vector having distribution function $F_{X_{n},X_{n+1}}:=B_{n,n+1}(F_{n},F_{n+1})$ defined through Sklar’s theorem. For Borel sets $A\in\mathscr{B}(\mathbb{R}^{n+1})$ and $B\in\mathscr{B}(\mathbb{R})$, we define the distribution $\mu_{\mathcal{T}}$ on $\mathcal{B}(\mathbb{R}^{n+2})$ by